# Learning Content-enhanced Mask Transformer for Domain Generalized Urban-Scene Segmentation

### v1

@@ 3.4 Network Architecture and Implementation Details
The image decoder from [6] uses the off-the-shelf multi-scale deformable attention Transformer (MSDeformAttn) [68] with the default setting in [68, 6]. By considering the image features from the Swin-Based encoder as input, every 6 MSDeformAttn layers are used to progressively up-sample the image features in ×$\times 32$, ×$\times 16$, ×$\times 8$, and ×$\times 4$, respectively. The $1/4$ resolution feature map is fused with the features from the Transformer decoder for dense prediction.
@@
Following the default setting of MaskFormer [5] and Mask2Former [6], the final loss function $\mathcal{L}$ is a linear combination of the binary cross-entropy loss $\mathcal{L}_{ce}$, dice loss $\mathcal{L}_{dice}$, and the classification loss $\mathcal{L}_{cls}$, given by
@@
4 Experiment
4.1 Dataset & Evaluation Protocols
Building upon prior research in domain-generalized USSS, our experiments utilize five different semantic segmentation datasets. Specifically, CityScapes [8] provides 2,975 and 500 well-annotated samples for training and validation, respectively. These driving-scenes are captured in Germany cities with a resolution of 2048$\times$1024. BDD-100K [59] also provides diverse urban driving scenes with a resolution of 1280$\times$720. 7,000 and 1,000 fine-annotated samples are provided for training and validation of semantic segmentation, respectively. Mapillary [33] is also a real-scene large-scale semantic segmentation dataset with 25,000 samples. SYNTHIA [43] is large-scale synthetic dataset, and provides 9,400 images with a resolution of 1280$\times$760. GTA5 [42] is a synthetic semantic segmentation dataset rendered by the GTAV game engine. It provides 24,966 simulated urban-street samples with a resolution of 1914$\times$1052. We use C, B, M, S and G to denote these five datasets.
Following prior domain generalized USSS works [35, 36, 7, 38], the segmentation model is trained on one dataset as the source domain, and is validated on the rest of the four datasets as the target domains. Three settings include: 1) G to C, B, M, S; 2) S to C, B, M, G; and 3) C to B, M, G, S. mIoU (in percentage %) is used as the validation metric. All of our experiments are performed three times and averaged for fair comparison. All the reported performance is directly cited from prior works under the ResNet-50 backbone [35, 36, 7, 38].
4.2 Comparison with State-of-the-art Domain Generalized USSS Methods
@@
Method Proc. & Year Trained on GTA5 (G) Trained on SYNTHIA (S) Trained on Cityscapes (C)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ S $\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
IBN [35] ECCV2018 33.85 32.30 37.75 27.90 32.04 30.57 32.16 26.90 48.56 57.04 45.06 26.14
IW [36] CVPR2019 29.91 27.48 29.71 27.61 28.16 27.12 26.31 26.51 48.49 55.82 44.87 26.10
Iternorm [17] CVPR2019 31.81 32.70 33.88 27.07 - - - - 49.23 56.26 45.73 25.98
DRPC [62] ICCV2019 37.42 32.14 34.12 28.06 35.65 31.53 32.74 28.75 49.86 56.34 45.62 26.58
ISW [7] CVPR2021 36.58 35.20 40.33 28.30 35.83 31.62 30.84 27.68 50.73 58.64 45.00 26.20
GTR [37] TIP2021 37.53 33.75 34.52 28.17 36.84 32.02 32.89 28.02 50.75 57.16 45.79 26.47
DIRL [57] AAAI2022 41.04 39.15 41.60 - - - - - 51.80 - 46.52 26.50
SHADE [65] ECCV2022 44.65 39.28 43.34 - - - - - 50.95 60.67 48.61 27.62
SAW [38] CVPR2022 39.75 37.34 41.86 30.79 38.92 35.24 34.52 29.16 52.95 59.81 47.28 28.32
WildNet [23] CVPR2022 44.62 38.42 46.09 31.34 - - - - 50.94 58.79 47.01 27.95
AdvStyle [66] NeurIPS2022 39.62 35.54 37.00 - 37.59 27.45 31.76 - - - - -
CMFormer (Ours) 2023 55.31 49.91 60.09 43.80 44.59 33.44 43.25 40.65 59.27 71.10 58.11 40.43
$\uparrow$10.66 $\uparrow$10.63 $\uparrow$14.00 $\uparrow$12.46 $\uparrow$5.67 $\downarrow$1.80 $\uparrow$8.73 $\uparrow$11.49 $\uparrow$6.32 $\uparrow$10.43 $\uparrow$ 9.50 $\uparrow$12.11
GTA5 Source Domain The third column of Table 1 reports the performance on target domains of C, B, M and S, respectively. The proposed CMFormer shows a performance improvement of 10.66%, 10.63%, 14.00% and 12.46% compared to existing state-of-the-art CNN based methods on each target domain, respectively. These outcomes demonstrate the feature generalization ability of the proposed CMFormer. Notice that the source domain GTA5 is a synthetic dataset, while the target domains are real images. It further validates the performance of the proposed method.
SYNTHIA Source Domain The forth column of Table 1 reports the performance. The proposed CMFormer shows a 5.67%, 8.73% and 11.49% mIoU performance gain against the best CNN based methods, respectively. However, on the BBD-100K (B) dataset, the semantic-aware whitening (SAW) method [38] outperforms the proposed CMFormer by 1.80% mIoU. Nevertheless, the proposed CMFormer still outperforms the rest state-of-the-art methods. The performance gain of the proposed CMFormer when trained on SYNTHIA dataset is not as significant as it is trained on CityScapes or GTA5 dataset. The explanation may be that the SYNTHIA dataset has much fewer samples than GTA5 dataset, i.e., 9400 v.s. 24966, and a transformer may be under-trained.
CityScapes Source Domain The last column of Table 1 reports the performance. The proposed CMFormer shows a performance gain of 6.32%, 10.43%, 9.50% and 12.11% mIoU on the B, M, G and S dataset against the state-of-the-art CNN based method. As BDD100K dataset contains many nigh-time urban-street images, it is particularly challenging for existing domain generalized USSS methods. Still, a performance gain of 6.32% is observed by the proposed CMFormer.
4.3 Ablation Studies
@@
Table 2: Ablation studies on each component of the proposed CMFormer. $\times$32, $\times$16 and $\times$8 denote the image features of $\times$32, $\times$16 and $\times$8 resolution. $\checkmark$ refers to the content enhancement is implemented. Evaluation metric mIoU is presented in (%).
Table 3: Ablation studies on the embedding order of the multi-resolution features in the CMFormer. Ours are compared with the plain and descent settings. Evaluation metric mIoU in %.
@@
Method Trained on Cityscapes (C)
$\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
$plain$ 59.20 71.08 57.89 40.16
$descent$ 58.51 71.01 57.53 40.29
Ours 59.27 71.10 58.11 40.43
Table 3: Ablation studies on the embedding order of the multi-resolution features in the CMFormer. Ours are compared with the plain and descent settings. Evaluation metric mIoU in %.
To better understand the content enhancement for mask attention, some segmentation masks learnt by the existing mask attention and our content enhanced mask attention are visualized in the second and third row of Fig. 4. The proposed content enhanced mask attention is less sensitive to the style variation. The learnt masks are more capable to separate the key objects from the background.
Figure 4: Visualization of the segmentation masks learnt by the original mask attention (in second row) and the proposed content-enhanced mask attention (in third row), under the C [8] $\rightarrow$ B [59], G [42], M [33], S [43] settings (in first row). The proposed content-enhanced mask attention is less sensitive to the style variation and is of better capability to separate the objects from background.
On Multi-resolution Feature Embedding Strategy is inherited from Mask2Former in an alternative manner [6]. To investigate its effectiveness, the method is compared with two alternative solutions: (1) plain sequence embedding (denoted as $plain$) and (2) alternative embedding in a descent order (denoted as $descent$). For the plain sequence embedding, the first, second and third three decoder layers are fed into the ×$\times 32$, ×$\times 16$ and ×$\times 8$ image features, respectively. For the alternative embedding in a descent order, each three decoder layers are fed into the image features with the order of ×$\times 8$, ×$\times 16$ and ×$\times 32$. The results are listed in Table 3. The embedding strategy inherited from [6] has a slightly better performance compared to the two other solutions.
4.4 Comparison with Recent State-of-the-art in Related Tasks
The proposed CMFormer is compared against three categories of recent state-of-the-art approaches that are relevant to related tasks. While these approaches do not specifically focus on domain-generalized USSS, we compare them to further highlight the optimality of the proposed CMFormer.
Specially, we consider mask-level Transformer segmentation models (Segmenter [47], MaskFormer [5], Mask2Former [6], denoted as $\mathcal{M}$), combining state-of-the-art CNN based domain generalization techniques with segmentation Transformer (denoted as $\mathcal{G}$), and the plain segmentation Transformer that is pre-trained from the foundation models (SAM [21], SegGPT [54], denoted as $\mathcal{F}$). It can be seen that the proposed CMFormer significantly outperforms other methods. It is also very interesting to observe that the whitening transformation (IW [36], ISW [36], SAW [38]) for CNN based domain generalized segmentation frameworks do not work for mask-level Transformer segmentation pipeline.
[…]
Method Backbone Category Trained on Cityscapes (C)
$\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
Segmenter [47] ViT-B $\mathcal{M}$ 31.42 37.76 28.56 17.32
Max-DeepLab [51] MaX-L 47.82 54.05 39.82 27.15
CMT-DeepLab [60] Axial-R104 48.74 57.93 41.66 26.06
kMaX-DeepLab [61] ConvNeXt-B 50.58 59.06 43.92 28.70
MaskFormer [5] Swin-B 51.86 64.70 49.24 33.35
Mask2Former [6] Swin-B 55.43 66.12 55.05 38.19
IN [35] + Mask2Former [6] Swin-B $\mathcal{G}$ 56.78 67.97 56.07 40.16
IW [36] + Mask2Former [6] Swin-B 54.20 63.14 50.57 35.59
ISW [36] + Mask2Former [6] Swin-B 53.00 63.16 52.14 36.97
SAW [38] + Mask2Former [6] Swin-B 55.78 62.36 57.88 39.96
SAM Pre-trained [21] ViT-H $\mathcal{F}$ 52.36 65.72 51.49 34.25
SegGPT Pre-trained [54] ViT-L 50.94 61.37 44.85 32.36
CMFormer (Ours) Swin-B $\mathcal{G}$ 59.27 71.10 58.11 40.43
4.5 Extra Experiments on Generalization
To Adverse Domains Beyond the standard validation protocols in existing methods [35, 16, 36, 7, 38, 66], we further validate the proposed CMFormer’s performance by benchmarking it on the adverse conditions dataset with correspondance (ACDC) [44]. It is the largest semantic segmentation dataset under a variety of adverse conditions, including rain, fog, night and snow. We set the fog, night, rain and snow as four different unseen domains, and directly use the model pre-trained on CityScapes for inference. The results are shown in Table 6. It significantly outperforms existing domain generalized segmentation methods (IN [35], Iternorm [16], IW [36], ISW [7], ISSA [25]) by upto 10.3%, 0.5%, 11.6%, 11.1% on the fog, night, rain and snow domains, respectively. On the other hand, it shows a superior performance of 4.4%, 4.0% and 1.8% against the second best-performed method Mask2Former on the fog, rain and snow domain, respectively.
[…]
Method Backbone Trained on Cityscapes (C)
$\rightarrow$ Fog $\rightarrow$ Night $\rightarrow$ Rain $\rightarrow$ Snow mean
IN [35] ResNet-50 63.8 21.2 50.4 49.6 43.7
Iternorm [16] ResNet-50 63.3 23.8 50.1 49.9 45.3
IW [36] ResNet-50 62.4 21.8 52.4 47.6 46.6
ISW [7] ResNet-50 64.3 24.3 56.0 49.8 48.1
ISSA [25] MiT-B2 67.5 33.2 55.9 53.2 52.5
RefineNet [26] ResNet-101 46.4 29.0 52.6 43.3 43.7
DeepLabv3+ [3] ResNet-101 45.7 25.0 50.0 42.0 41.6
DANet [12] ResNet-101 34.7 19.1 41.5 33.3 33.1
HRNet [52] HR-w48 38.4 20.6 44.8 35.1 35.3
SegFormer [56] MiT-B2 59.2 38.9 62.5 58.2 56.2
Mask2Former [6] Swin-B 73.4 37.1 63.6 62.5 58.0
CMFormer (Ours) Swin-B 77.8 33.7 67.6 64.3 60.1
Backbone Trained on Synthetic Domain (G+S)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M mean
ResNet-50 35.46 25.09 31.94 30.83
IBN [35] 35.55 32.18 38.09 35.27
ISW [7] 37.69 34.09 38.49 36.75
SHADE [65] 47.43 40.30 47.60 45.11
AdvStyle [66] 39.29 39.26 41.14 39.90
Ours 59.70 53.36 61.61 58.22
[…]
4.6 Visualization
T-SNE visualization To better understand the style change handling and stable robust representation learning, the feature space with samples when using B, M, G, S as target domains are shown in Fig. 5 by t-SNE. ISW [7], a typical style-decoupling method, has some difficulty to handle the style variation, but the distance of the semantic embeddings from each domain is close. Advstyle [66], a typical adverse style augmentation method, can better mix the styles, but the distance of semantic embeddings is not properly reduced. The original mask attention from [6] can reduce the distance of each semantic embedding but still has some difficulty to mix the different styles. In contrast, the proposed CMFormer can not only reduce the distance of semantic embeddings, but also allow the styles to be more uniformly distributed and therefore minimizes the domain gap.
Figure 5: T-SNE visualization of the features space on target domains. (a) ISW [7], (b) AdvStyle [66], (c) Mask2Former [6], (d) the proposed CMFormer. Our method has the most uniform distribution, which means the domain gap is minimized.
@@ 4.6 Visualization
Quantitative segmentation prediction Some generalized segmentation prediction results on the C $\rightarrow$ B, M, G, S setting (in Fig. 9) and on the C $\rightarrow$ adverse domain setting (in Fig. 10). Compared with the CNN based methods, the proposed CMFormer shows a better segmentation prediction, especially in terms of the completeness of objects.
[…]
Figure 7: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN [35], IW [36], Iternorm [17], ISW [7]) and the proposed CMFormer under the C $\rightarrow$ adverse domain setting. More prediction is in the supplementary material.
5 Conclusion
@@
In this paper, we explored the feasibility of adapting the mask Transformer for domain-generalized urban-scene semantic segmentation (USSS). To address the challenges of style variation and robust content representation, we proposed a content-enhanced mask attention mechanism. This mechanism is designed to capture more resilient content features while being less sensitive to style variations. Furthermore, we extended it to incorporate multi-resolution features and integrate it into a novel framework called the Content-enhanced Mask TransFormer (CMFormer). To evaluate the effectiveness of CMFormer, we conducted extensive experiments on multiple settings. The results demonstrated the superior performance of CMFormer compared to existing domain-generalized USSS methods.
Limitation Discussion & Boarder Social Impact. The proposed content-enhanced mask attention mechanism is derived from the self-attention mechanism and can be seamlessly integrated into ViT based segmentation pipelines. The experimental results demonstrate its superior performance, thereby indicating the potential to shift the focus of domain-generalized USSS towards ViT based pipelines.
@@
Given the criticality of safety-related road applications, the proposed method holds significant importance. It has the potential to enhance the accuracy and reliability of semantic segmentation models, thereby contributing to safer and more efficient autonomous systems. Overall, the proposed content-enhanced mask attention mechanism not only offers promising advancements in domain-generalized USSS but also holds potential for broader applications in real-world scenarios.

### v4

@@ Introduction
The image decoder from (Cheng et al. 2022) uses the off-the-shelf multi-scale deformable attention Transformer (MSDeformAttn) (Zhu et al. 2021) with the default setting in (Zhu et al. 2021; Cheng et al. 2022). By considering the image features from the Swin-Based encoder as input, every 6 MSDeformAttn layers are used to progressively up-sample the image features in ×$\times 32$, ×$\times 16$, ×$\times 8$, and ×$\times 4$, respectively. The $1/4$ resolution feature map is fused with the features from the Transformer decoder for dense prediction.
The Transformer decoder is also directly inherited from Mask2Former (Cheng et al. 2022), which has 9 self-attention layers in the Transformer decoder to handle the ×$\times 32$, ×$\times 16$ and ×$\times 8$ image features, respectively.
@@
Following the default setting of MaskFormer (Cheng, Schwing, and Kirillov 2021) and Mask2Former (Cheng et al. 2022), the final loss function $\mathcal{L}$ is a linear combination of the binary cross-entropy loss $\mathcal{L}_{ce}$, dice loss $\mathcal{L}_{dice}$, and the classification loss $\mathcal{L}_{cls}$, given by
@@
Experiment
Dataset & Evaluation Protocols
Building upon prior research in domain-generalized USSS, our experiments utilize five different semantic segmentation datasets. Specifically, CityScapes (Cordts et al. 2016) provides 2,975 and 500 well-annotated samples for training and validation, respectively. These driving-scenes are captured in Germany cities with a resolution of 2048$\times$1024. BDD-100K (Yu et al. 2018) also provides diverse urban driving scenes with a resolution of 1280$\times$720. 7,000 and 1,000 fine-annotated samples are provided for training and validation of semantic segmentation, respectively. Mapillary (Neuhold et al. 2017) is also a real-scene large-scale semantic segmentation dataset with 25,000 samples. SYNTHIA (Ros et al. 2016) is large-scale synthetic dataset, and provides 9,400 images with a resolution of 1280$\times$760. GTA5 (Richter et al. 2016) is a synthetic semantic segmentation dataset rendered by the GTAV game engine. It provides 24,966 simulated urban-street samples with a resolution of 1914$\times$1052. We use C, B, M, S and G to denote these five datasets.
Following prior domain generalized USSS works (Pan et al. 2018, 2019; Choi et al. 2021; Peng et al. 2022), the segmentation model is trained on one dataset as the source domain, and is validated on the rest of the four datasets as the target domains. Three settings include: 1) G to C, B, M, S; 2) S to C, B, M, G; and 3) C to B, M, G, S. mIoU (in percentage %) is used as the validation metric. All of our experiments are performed three times and averaged for fair comparison. All the reported performance is directly cited from prior works under the ResNet-50 backbone (Pan et al. 2018, 2019; Choi et al. 2021; Peng et al. 2022).
Existing domain generalized USSS methods are included for comparison, namely, IBN (Pan et al. 2018), IW (Pan et al. 2019), Iternorm (Huang et al. 2019b), DRPC (Yue et al. 2019b), ISW (Choi et al. 2021), GTR (Peng et al. 2021), DIRL (Xu et al. 2022), SHADE (Zhao et al. 2022), SAW (Peng et al. 2022), WildNet (Lee et al. 2022), AdvStyle (Zhong et al. 2022), SPC (Huang et al. 2023), HGFormer (Ding et al. 2023).
Comparison with State-of-the-art
Method Proc. & Year Trained on GTA5 (G)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ S
IBN ECCV2018 33.85 32.30 37.75 27.90
IW CVPR2019 29.91 27.48 29.71 27.61
Iternorm CVPR2019 31.81 32.70 33.88 27.07
DRPC ICCV2019 37.42 32.14 34.12 28.06
ISW CVPR2021 36.58 35.20 40.33 28.30
GTR TIP2021 37.53 33.75 34.52 28.17
DIRL AAAI2022 41.04 39.15 41.60 -
SHADE ECCV2022 44.65 39.28 43.34 -
SAW CVPR2022 39.75 37.34 41.86 30.79
WildNet CVPR2022 44.62 38.42 46.09 31.34
AdvStyle NeurIPS2022 39.62 35.54 37.00 -
SPC CVPR2023 44.10 40.46 45.51 -
HGFormer CVPR2023 - - - -
CMFormer (Ours) 2023 55.31 49.91 60.09 43.80
$\uparrow$10.66 $\uparrow$9.45 $\uparrow$14.00 $\uparrow$12.46
@@
GTA5 Source Domain Table 1 reports the performance on target domains of C, B, M and S, respectively. The proposed CMFormer shows a performance improvement of 10.66%, 9.45%, 14.00% and 12.46% compared to existing state-of-the-art CNN based methods on each target domain, respectively. These outcomes demonstrate the feature generalization ability of the proposed CMFormer. Notice that the source domain GTA5 is a synthetic dataset, while the target domains are real images. It further validates the performance of the proposed method.
SYNTHIA Source Domain Table 2 reports the performance. The proposed CMFormer shows a 5.67%, 8.73% and 11.49% mIoU performance gain against the best CNN based methods, respectively. However, on the BBD-100K (B) dataset, the semantic-aware whitening (SAW) method (Peng et al. 2022) outperforms the proposed CMFormer by 1.80% mIoU. Nevertheless, the proposed CMFormer still outperforms the rest state-of-the-art methods. The performance gain of the proposed CMFormer when trained on SYNTHIA dataset is not as significant as it is trained on CityScapes or GTA5 dataset. The explanation may be that the SYNTHIA dataset has much fewer samples than GTA5 dataset, i.e., 9400 v.s. 24966, and a transformer may be under-trained.
Method Proc. & Year Trained on SYNTHIA (S)
$\rightarrow$ C $\rightarrow$ B $\rightarrow$ M $\rightarrow$ G
IBN ECCV2018 32.04 30.57 32.16 26.90
IW CVPR2019 28.16 27.12 26.31 26.51
Iternorm CVPR2019 - - - -
DRPC ICCV2019 35.65 31.53 32.74 28.75
ISW CVPR2021 35.83 31.62 30.84 27.68
GTR TIP2021 36.84 32.02 32.89 28.02
DIRL AAAI2022 - - - -
SHADE ECCV2022 - - - -
SAW CVPR2022 38.92 35.24 34.52 29.16
WildNet CVPR2022 - - - -
AdvStyle NeurIPS2022 37.59 27.45 31.76 -
SPC CVPR2023 - - - -
HGFormer CVPR2023 - - - -
CMFormer (Ours) 2023 44.59 33.44 43.25 40.65
$\uparrow$5.67 $\downarrow$1.80 $\uparrow$8.73 $\uparrow$11.49
@@
CityScapes Source Domain Table 3 reports the performance. As HGFormer only reports one decimal results (Ding et al. 2023), we also report one decimal results when compared with them. The proposed CMFormer (with Swin-Base backbone) shows a performance gain of 6.32%, 10.43%, 9.50% and 12.11% mIoU on the B, M, G and S dataset against the state-of-the-art CNN based method. As BDD100K dataset contains many nigh-time urban-street images, it is particularly challenging for existing domain generalized USSS methods. Still, a performance gain of 6.32% is observed by the proposed CMFormer.
On the other hand, when comparing ours with the contemporary HGFormer with the Swin-Large backbone, it shows an mIoU improvement of 1.1%, 1.5%, 1.3% and 1.7% on the B, M, G and S target domain, respectively.
Method Backbone Trained on Cityscapes (C)
$\rightarrow$ B $\rightarrow$ M $\rightarrow$ G $\rightarrow$ S
IBN ResNet50 48.56 57.04 45.06 26.14
IW ResNet50 48.49 55.82 44.87 26.10
Iternorm ResNet50 49.23 56.26 45.73 25.98
DRPC ResNet50 49.86 56.34 45.62 26.58
ISW ResNet50 50.73 58.64 45.00 26.20
GTR ResNet50 50.75 57.16 45.79 26.47
DIRL ResNet50 51.80 - 46.52 26.50
SHADE ResNet50 50.95 60.67 48.61 27.62
SAW ResNet50 52.95 59.81 47.28 28.32
WildNet ResNet50 50.94 58.79 47.01 27.95
AdvStyle ResNet50 - - - -
SPC ResNet50 - - - -
[…]
Mask2Former Swin-B 55.43 66.12 55.05 38.19
CMFormer (Ours) Swin-B 59.27 71.10 58.11 40.43
[…]
CMFormer (Ours) Swin-L 62.6 73.6 60.7 43.0
[…]
Backbone Trained on Two Synthetic Domains (G+S)
$\rightarrow$ Citys $\rightarrow$ BDD $\rightarrow$ MAP mean
ResNet-50 35.46 25.09 31.94 30.83
IBN 35.55 32.18 38.09 35.27
ISW 37.69 34.09 38.49 36.75
SHADE 47.43 40.30 47.60 45.11
AdvStyle 39.29 39.26 41.14 39.90
SPC 46.36 43.18 48.23 45.92
Ours 59.70 53.36 61.61 58.22
[…]
From Clear to Adverse Conditions The results and discussions are provided in the supplementary material.
[…]
Figure 6: Unseen domain segmentation prediction of existing CNN based domain generalized semantic segmentation methods (IBN (Pan et al. 2018), IW (Pan et al. 2019), ISW (Choi et al. 2021), SAW (Peng et al. 2022)) and the proposed CMFormer under the C $\rightarrow$ adverse domain setting. More prediction is in the supplementary material.
Ablation Studies
[…]
@@
Table 5: Ablation studies on each component of the proposed CMFormer. $\times$32, $\times$16 and $\times$8 denote the image features of $\times$32, $\times$16 and $\times$8 resolution. $\checkmark$ refers to the content enhancement is implemented. Evaluation metric mIoU.
Quantitative Segmentation Results
@@
Some generalized segmentation results on the C $\rightarrow$ B, M, G, S setting (in Fig. 5) and on the C $\rightarrow$ adverse domain setting (in Fig. 6). Compared with the CNN based methods, the proposed CMFormer shows a better segmentation prediction, especially in terms of the completeness of objects.
Conclusion
@@
In this paper, we explored the feasibility of adapting the mask Transformer for domain-generalized urban-scene semantic segmentation (USSS). To address the challenges of style variation and robust content representation, we proposed a content-enhanced mask attention (CMA) mechanism. This mechanism is designed to capture more resilient content features while being less sensitive to style variations. Furthermore, we extended it to incorporate multi-resolution features and integrate it into a novel framework called the Content-enhanced Mask TransFormer (CMFormer). To evaluate the effectiveness of CMFormer, we conducted extensive experiments on multiple settings. The results demonstrated the superior performance of CMFormer compared to existing domain-generalized USSS methods.
@@
Boarder Social Impact. Given the criticality of safety-related road applications, the proposed method holds significant importance. It has the potential to enhance the accuracy and reliability of semantic segmentation models, thereby contributing to safer and more efficient autonomous systems. Overall, the proposed content-enhanced mask attention mechanism not only offers promising advancements in domain-generalized USSS but also holds potential for broader applications in real-world scenarios.